OpenStudy (anonymous):

a cylindrical tank of height equal to twice the diameter of its base can hold 10,000 cm^3 of water. another cylindrical container with the same capacity has its height equal to 3 times the diameter of its base. find the ratio of the amount of aluminum required for making the two containers including covers

OpenStudy (campbell_st):

well look at the volume 1st and use the information to find the diameter tank 1 d = diameter r = d/2 then height = 2d or 4r so the volume is \[v = \pi \times r^2 \times 4r\] or \[10000 = 4\pi r^3\] solve for r tank 2 d = diameter then r = d/2 height = 3d or 6r use the volume information to find d \[10000 = \pi \times r^2 \times 6r\] or \[10000 = 6\pi r^3\] again solve for r. so now you have the diameter of each tank now the surface area tank 1 \[SA = 2\pi r^2 + 2\pi r \times 4r\] substitute the value of r you found from the volume of tank 1. tank 2 \[SA = 2\pi r^2 + 2\pi r \times 3r\] substitute the value of r you found from take 2. then make the ratio of surface areas tank 1 :tank 2 hope it helps